OpenStudy (wcrmelissa2001):

Find the value(s) of x for which f(x)=f^-1(x) where f(x)=sqrt(x+2)-2/3 f^-1(x)=(x+2/3)^2-2

OpenStudy (wcrmelissa2001):

er I know f-1f(x)=x

OpenStudy (wcrmelissa2001):

as in I know how to get f^-1 but I don't know how to go on from there

OpenStudy (priyar):

oops sorry i made an error..! forget the abv statement..!

OpenStudy (priyar):

oh they have given the value of both... so just equate them!

OpenStudy (priyar):

\[\sqrt{x+2}-2/3 = (x+2/3) ^2 - 2\] ok?

OpenStudy (wcrmelissa2001):

yup but I'm not sure how to equate. That's the problem

OpenStudy (wcrmelissa2001):

As in I end up with x^4 or something like that and I'm not sure how to solve from there

OpenStudy (priyar):

yeah i see your problem..! then that is not the way to do it.. a easier way is to use the abv property that you wrote..

OpenStudy (priyar):

\[f(f^{-1}(x)) = x\] right? how can you rewrite this here..?

OpenStudy (priyar):

@wcrmelissa2001 ?

OpenStudy (wcrmelissa2001):

sorry im here

OpenStudy (wcrmelissa2001):

wait but if you were to do that then how would you explain that they are both equal?

OpenStudy (priyar):

yes use this info in the above eq..

OpenStudy (wcrmelissa2001):

no, as in I don't understand how the f-1f(x)=x will work in this case since it doesn't state that f-1(x)=f(x)???

OpenStudy (priyar):

no i meant more like:\[f(f(x)) =x\]..but even here there is a term with x^4..hmm.. let me see if i can think of a better method..otherwise we must do it the long way..

OpenStudy (wcrmelissa2001):

I tried drawing two graphs then finding their intersection too but nope

OpenStudy (phi):

inverse functions are symmetric about the line y=x if they are ever equal, it is when they meet on that line. so for f(x) = sqrt(x+2)-⅔ i.e. y = sqrt(x+2) - ⅔ we can replace the y with x to find the spot on the y=x line where the function crosses \[ x = \sqrt{x+2} - \frac{2}{3} \]

OpenStudy (wcrmelissa2001):

but that is the original function?

OpenStudy (phi):

maybe I have to rethink this. But here is a graph from wolfram http://www.wolframalpha.com/input/?i=plot+y%3D(x%2B%E2%85%94)%5E2+-2;+y%3D+sqrt(x%2B2)+-+%E2%85%94;+y%3D+x+for+x%3D-2+to+%2B2

OpenStudy (phi):

this plot shows the symmetry better, but my idea only gets one of the solutions

OpenStudy (wcrmelissa2001):

i plotted the graph too :(

OpenStudy (phi):

what course is this ? what grade ?

OpenStudy (wcrmelissa2001):

not sure... I don't take grades. Different system :D

OpenStudy (phi):

how much math are you supposed to know ?

OpenStudy (phi):

The brute force way is solve the quartic , which is possible , though painful by hand but is there an elegant way ? Off hand I don't know. (though we can find the positive solution easily enough) (-1+ sqrt(57))/6

OpenStudy (zarkon):

\[f(x)=\sqrt{x+2}-\frac{2}{3}\] therefore \(f:[-2,\infty)\to[-2/3,\infty)\) hence \[f^{-1}(x)=\left(x+\frac{2}{3}\right)^2-2\] and \(f^{-1}:[-2/3,\infty)\to[-2,\infty)\)

OpenStudy (wcrmelissa2001):

sorry i'm not sure i understand your working?

OpenStudy (loser66):

To me, just let them equal and solve for x. It is a mess but doable still. At the end up, you can have at most 4 values of x.

OpenStudy (phi):

Let g(x) be the inverse of f(x). Then we want to solve f(x) = g(x) This gives a 4th order equation. Messy to solve unless you use wolfram or mathematica. Here is one way to solve the quartic equation, in "chunks" the idea is the 4 roots of the quartic means you can factor it into (x-a)(x-b)(x-c)(x-d) = 0 using the idea that two of the roots lie on the line y=x we can solve (either) equation: f(x) = x this is a quadratic and gives two of the roots. One of them is what we want. to find the other two roots, we divide the quartic that arises from f(x)=g(x) by the quadratic that arises from f(x)=x this gives us a new quadratic. in other words, if we found two roots a, and b we are dividing the quartic by (x-a)(x-b) to get (x-c)(x-d) which we can now solve using the quadratic formula. Still a bit of a mess, but doable.

OpenStudy (zarkon):

I think everyone here needs to read the definition of an inverse function again https://en.wikipedia.org/wiki/Inverse_function#Definitions

OpenStudy (wcrmelissa2001):

Thanks everyone! :D Will go and try it out and keep trying and trying :( HAHA